# VibES: Induced Vibration for Persistent Event-based Sensing

### v1

@@ 4.3 Image Quality Evaluation
Scene Status Entropy $\uparrow$
@@
Scene Status NIQE $\downarrow$
@@
VIBES $\mathbf{5.1}\pm 0.3$
Pattern
Checkerboard
S-EV $8.2\pm 9.0$
VIBES $\mathbf{6.3}\pm 0.2$
@@ 4.4 Edge Extraction Evaluation
Scene Status Variance $\uparrow$ Grad. Mag. $\uparrow$
@@
Scene Status Avg. Num. Components $\downarrow$ Avg. Contour Length $\uparrow$ Avg. Junction Count
@@
V-EV $\mathbf{175.0\pm 52.5}$ $\mathbf{32.1\pm 10.8}$ $979.3\pm 349.2$
Pattern
Checkerboard
S-EV $300.3\pm 232.9$ $6.7\pm 5.3$ $193.2\pm 176.3$
V-EV $\mathbf{140.7\pm 40.2}$ $\mathbf{49.9\pm 17.3}$ $691.3\pm 63.2$
@@ 4.5 Runtime Evaluation
Our software stack, implemented in C++ with the Metavision API by Prophesee, is evaluated by running the compensation script ten times per dataset. The average processing time is $15.28\pm 3.7$ ns per event ($\sim 65$ Mev/s), including undistortion and motion compensation. The initial parameters estimation Sec. 3.2.2, requires $104.8\pm 0.03$ ms, but remains non-blocking thanks to a multithreaded design that tracks incoming events while buffering and processing compensation. A video demonstration is provided in the supplementary material. Although runtime is fast, real-time performance depends on sensor resolution and scene dynamics; in our experiments, event rates peaked at $45$ Mev/s.
@@ 5.1 Frequency Estimation
While our software stack is designed to detect the frequency of the camera’s oscillatory motion, it can also be applied to estimate the vibration frequency of an object in the scene with a static camera.
@@
The results in Tab. 5 demonstrate that VIBES achieves accurate frequency estimation with low bias and variance. We note, however, a slight increase in bias at higher frequencies, with deviations of up to 0.1Hz between the ground truth and the predicted values. This effect is likely due to aliasing, as we are able to estimate higher frequencies both in the simulated Sec. 5.2 and in real-world data.
@@ 5.2 Relative Depth Awareness
We explore how parallax induced by the camera vibration can encode information about scene depth.
Note that this type of depth estimation is not possible in setups like the wedge-prism solution proposed by [24], because the camera center of projection does not change.
We tested three synthetic scenes containing two staggered patterns at different depth planes, as in Fig. 6. To create synthetic data, we generated RGB frames using Blender [8] at 1000 FPS and converted them to events using V2E [9]. Camera motion was simulated according to the sinusoidal motion model described in Sec. 3.1. We instantiated two trackers per pattern. Then we stored the amplitudes estimated by the EKF for each of the patterns and averaged them across the entire scene duration. Finally, we computed the ratio between the average amplitudes to get the estimation of the relative depth. We repeated this process ten times per scene. Our simulated results show that our system reliably predicts the relative depth of the two patterns: for ground-truth relative depth ratios of $0.33$, $0.50$, and $0.66$ the predicted ratios are $0.41\pm 0.01$, $0.59\pm 0.02$, and $0.80\pm 0.03$. Please see Supplementary Sec. S2.6 for additional details.
Figure 6: Illustration of depth prediction setup. The two patterns are placed at different distances from the camera and we are capable of estimating the ratio $\frac{d_{1}}{d_{2}}$ (or vice versa) by observing the amplitudes of the perceived motion of the camera measured by the tracker (Sec. 3.2.1).
@@ 6 Conclusion
In this work, we presented VIBES, a novel bio-inspired method that enhances event cameras by actively inducing motion. Our approach produces persistent event streams for static scenes, overcoming a common obstacle in event-based vision.
We see several promising avenues for future work. For example, the vibration system could be made adaptive — i.e., disabled in highly dynamic scenes with dense event streams, and reactivated when the event density drops. Such an approach would preserve the low-power benefits of event cameras while ensuring a continuous visual signal when needed. Additionally, although our method already handles high event rates, downsampling or other event filtering techniques may be considered to further improve the real time performance, given the high event bandwidth generated by the induced motion. Our approach could also be combined with other event-based algorithms to improve performance on downstream tasks.
@@
Figure S3: A picture of the real prototype
@@
Shannon Entropy [40]: We calculate Shannon Entropy over binary accumulated event images, where the binary pixels are treated as drawn from the same distribution. This is calculated as follows:
@@
Natural Image Quality Evaluator (NIQE) [33]:
@@
Variance of image pixel values: Following [18] which found that high variance in accumulated event frames corresponds to sharper edges. We generated accumulated frames over a 10 ms window, and then calculated the variance of the per-pixel event counts:
@@
Gradient magnitude of image pixel values: We perform the same event accumulation procedure as above, but calculate per-pixel gradients and take an average of the magnitude over the image. We approximate per-pixel gradients using forward differences which can be calculated with convolutions:
@@
Edges continuity and fragmentation:
Once the raw edge maps are smoothed and thinned, we calculate three metrics.
@@
S2.6 Details on Camera Distance and Correction
@@
Methodology Motor Type
Power
Consumption
$\downarrow$
Measurement/Evaluation Basis
@@
Testa et al. [43] PTU-D46-17 [16] 1W Holding Power Off
6W Low Power
@@
Testa et al. [44] PTU-E46 [17] 1W Holding Power Off
6W Low Power
@@
We report the power consumption of our rotating unbalanced mass system and compare its performance against alternative approaches. VIBES employs a DC hobby motor, with parameters listed in Sec. S1.1.1. From our experiments, we observe that the motor draws an average current of 0.1410.141A, while the applied voltage was set at 22V. As most trials were conducted at 22V, we report the typical power consumption of our DC motor setup as
@@
Power consumption values are summarized in Tab. S3. For each methodology, we report both the estimated power consumption and the basis of evaluation. When specific hardware is identified, we reference the corresponding manufacturer’s datasheets. For example, the FLIR PTU units [16, 17] list distinct power requirements for holding (power off), low-power, and full-power modes. For VIBES, power consumption was measured directly during online experiments. For other approaches, expected values are reported, defaulting to no-load or standby power consumption when both online measurements and datasheet values are unavailable.

### v2

@@ 4.3 Image Quality Evaluation
Scene Method Entropy $\uparrow$
@@
Scene Method NIQE $\downarrow$
@@
V-EV $\mathbf{5.1}\pm 0.3$
Pattern Checkerboard S-EV $8.2\pm 9.0$
V-EV $\mathbf{6.3}\pm 0.2$
@@ 4.4 Edge Extraction Evaluation
Scene Method Variance $\uparrow$ Grad. Mag. $\uparrow$
@@
Scene Method Avg. Num. Components $\downarrow$ Avg. Contour Length $\uparrow$ Avg. Junction Count
@@
VibES $\mathbf{175.0\pm 52.5}$ $\mathbf{32.1\pm 10.8}$ $979.3\pm 349.2$
Pattern Checkerboard S-EV $300.3\pm 232.9$ $6.7\pm 5.3$ $193.2\pm 176.3$
VibES $\mathbf{140.7\pm 40.2}$ $\mathbf{49.9\pm 17.3}$ $691.3\pm 63.2$
@@ 4.5 Runtime Evaluation
Our software stack, implemented in C++ with the Metavision API by Prophesee, is evaluated by running the compensation script ten times per dataset. The average processing time is $15.28\pm 3.7$ ns per event ($\sim 65$ Mev/s), including undistortion and motion compensation. The initial parameter estimation (Sec. 3.2.2), requires $104.8\pm 0.03$ ms, but remains non-blocking thanks to a multithreaded design that tracks incoming events while buffering and performing motion compensation. A video demonstration is provided in the supplementary material. Although runtime is fast, real-time performance depends on sensor resolution and scene dynamics; in our experiments, event rates peaked at $45$ Mev/s.
@@ 5.1 Frequency Estimation
While our software is designed to detect the frequency of the camera’s oscillatory motion, it can also be applied to estimate the vibration frequency of an object in the scene with a static camera.
@@
The results in Tab. 5 demonstrate that VibES achieves accurate frequency estimation with low bias and variance. We note, however, a slight increase in bias at higher frequencies, with deviations of up to 0.1 Hz between the ground truth and the predicted values. This effect is likely due to aliasing, as we are able to estimate higher frequencies in both simulated data Sec. 5.2 and real-world data.
@@ 5.2 Relative Depth Awareness
We explore how parallax induced by the camera vibration can encode information about scene depth. Note that this type of depth estimation is not possible in setups like the wedge-prism solution proposed by [24], because the camera center of projection does not change.
We test three synthetic scenes containing two staggered patterns at different depth planes, as in Fig. 6. To create synthetic data, we generate RGB frames using Blender [8] at 1000 FPS and convert them to events using V2E [9]. Camera motion is simulated according to the sinusoidal motion model described in Sec. 3.1. We instantiate two trackers per pattern. Then we store the amplitudes estimated by the EKF for each of the patterns and average them across the entire scene duration. Finally, we compute the ratio between the average amplitudes to get the estimation of the relative depth. Our simulated results show that our system reliably predicts the relative depth of the two patterns: for ground-truth relative depth ratios of $0.33$, $0.50$, and $0.66$ the predicted ratios are $0.41\pm 0.01$, $0.59\pm 0.02$, and $0.80\pm 0.03$. Please see Supplementary Sec. S2.6 for additional details.
Figure 6: Illustration of depth prediction setup. The two patterns are placed at different distances from the camera and we estimate the ratio $\frac{d_{1}}{d_{2}}$ (or vice versa) by observing the amplitudes of the perceived motion of the camera measured by the tracker (Sec. 3.2.1).
@@ 6 Conclusion
In this work, we introduced VibES, a bio-inspired approach that enhances event cameras by actively inducing motion. Our method generates persistent event streams in static scenes, addressing a key limitation of event-based vision.
We see several promising avenues for future work. For example, the vibration system could be made adaptive—that is, disabled in highly dynamic scenes with dense event streams, and reactivated when the event density drops. Such an approach would preserve the low-power benefits of event cameras while ensuring a continuous visual signal when needed. Additionally, although our method already handles high event rates, downsampling or other event filtering techniques may be considered to further improve the real time performance, given the high event bandwidth generated by the induced motion. Our approach could also be combined with other event-based algorithms to improve performance on downstream tasks.
Acknowledgments
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through the RGPIN program and the Canada Research Chairs program, the Canada Foundation for Innovation, and the Ontario Research Fund.
@@ Acknowledgments
Figure S3: A photo of the hardware prototype.
@@
Shannon Entropy [40]. We calculate Shannon Entropy over binary accumulated event images, where the binary pixels are treated as drawn from the same distribution. This is calculated as follows:
@@
Natural Image Quality Evaluator (NIQE) [33].
@@
Variance of image pixel values. Following [18] which found that high variance in accumulated event frames corresponds to sharper edges. We generated accumulated frames over a 10 ms window, and then calculated the variance of the per-pixel event counts:
@@
Gradient magnitude of image pixel values. We perform the same event accumulation procedure as above, but calculate per-pixel gradients and take an average of the magnitude over the image. We approximate per-pixel gradients using forward differences which can be calculated with convolutions:
@@
Edges continuity and fragmentation. Once the raw edge maps are smoothed and thinned, we calculate three metrics.
@@
S2.6 Limits of Depth Estimation under Camera-Induced Motion
@@
Methodology Motor Type Power Consumption $\downarrow$ Measurement/Evaluation Basis
@@
Testa et al. [43] PTU-D46-17 [16] 6W Low Power
@@
Testa et al. [44] PTU-E46 [17] 6W Low Power
@@
We report the power consumption of our rotating unbalanced mass system and compare its performance against alternative approaches. VibES employs a DC hobby motor, with parameters listed in Sec. S1.1.1. From our experiments, we observe that the motor draws an average current of 0.1410.141A, while the applied voltage was set at 22V. Consequently, the typical power consumption of our DC motor setup is
@@
Power consumption values are summarized in Tab. S3. For each methodology, we report both the estimated power consumption and the basis of evaluation. When specific hardware is identified, we reference the corresponding manufacturer’s datasheets. For example, the FLIR PTU units [16, 17] list distinct power requirements for low-power and full-power modes. For VibES, power consumption was measured directly during online experiments. For other approaches, expected values are reported, defaulting to no-load or standby power consumption when both online measurements and datasheet values are unavailable.